Find f(x), assuming that f(x) ex dx = f(x) e - 8x-1 ex dx. (Use C for the constant of integration.)Evaluate the integral. (Use C for the constant of integration.) cos 498(3y) sin?(3y) dy
Answer #1

Solution: Given [f(x)e*dx =f(x)e* – 18x=tex dx de sfat)e*dx = tf(x)e) - 18x-te* dx dx [:(uv)=uv+uv] =>{{x}ef = f(x)e* +